# Charged-Current Elastic Scattering at the Electron-Ion Collider

### v1

@@ IV Veto System
As an example of a successful veto system, layered plastic scintillator detectors with single-layer efficiencies of 99.8% for single charged particles were built and operated for the NA62 experiment Ambrosino et al. (2016). Three layers of similar detectors arranged continuously around the beampipe in a configuration similar to the one sketched in Fig. 6 could provide the necessary veto efficiency of $\approx 10^{6}$. The shape of the scintillators will necessarily conform to the final shape of the beampipe, which has a complicated geometry in the far-forward region Aschenauer et al. (2025).
@@
A unique challenge for a scintillator-based charged particle veto system at the EIC is that the instantaneous rate of hits per channel can be large, depending on the granularity that is chosen. The expected averaged event rate at the EIC at top luminosity is around 500 kHz, and a large fraction of those events will produce hits in the forward region. A non-negligible hit rate from beam-gas events and synchrotron radiation can also be expected. The bunches at the EIC are spaced by roughly 10 nanoseconds. Two closely timed hits, e.g. from two consecutive bunch crossings, in the scintillator can result in the later hit not being properly vetoed if the response time of the scintillator and photosensor is too long. From that perspective, a photosensor with a fast rise- and fall-time such as an MCP-PMT may be desired to avoid hits shortly after one another piling up and defeating the veto. The decay time of plastic scintillator, for example formulas based on PPO fluorescence with POPOP wavelength shifter, is commonly as short at 2 ns, making it ideal for this use case Technology (nd). The expected radiation dose in the EIC far-forward region is approximately $3\times 10^{11}$ 1 MeV neutron equivalent of non-ionizing radiation and 100 krad of ionizing radiation in 100 fb-1. Plastic scintillators have been shown to retain their performance at significantly higher radiation doses Kharzheev (2019). Finally, plastic scintillator can be extruded or molded into awkward shapes, as will likely be necessary for hermetic coverage around the beampipe.
V Remarks on Charged Current Cross Sections
@@
It can be seen from Fig. 2 that the CCE total cross section remains approximately constant with $\sqrt{s}$ due to competition between the nucleon form factors and increased effective flux of $W^{-}$. On the other hand, the inclusive charged-current cross section increases linearly with $E_{\nu}$ due to increasing contribution of charged-current DIS, which dominates above $E_{\nu}\approx 20$ GeV. The phase space in $\gamma p$ center-of-mass energy $W$ for hard exclusive processes such as deeply virtual Compton scattering (DVCS) and deeply virtual meson production (DVMP) are much larger than for elastic scattering, and the DVCS and DVMP cross sections are known to grow with $W$ Aktas et al. (2005); Chekanov et al. (2009); Aaron et al. (2010); Kumerički and Mueller (2010). It may be the case then that at EIC energies, the cross sections for charged-current exclusive processes such as CCDVCS ($e^{-}p\rightarrow\nu_{e}n\gamma$) or CCDVMP (e.g. $e^{-}p\rightarrow\nu_{e}p\pi^{-}$) could be larger than the CCE cross section and counterintuitively may be more experimentally favorable. In particular, the $e^{-}p\rightarrow\nu_{e}p\pi^{-}$ reaction would have improved resolution on missing transverse momentum and missing mass since they could be measured in trackers. Theoretical estimates of these cross sections would therefore be interesting to see.
VI Conclusion
@@
The charged-current elastic scattering process provides a novel means to study the spin structure of the nucleon and potentially some presently unexplored corners of BSM physics. With its high luminosity and polarized beams of electrons and protons, the EIC is, in principle, an excellent tool for studying this reaction. However, the presence of a dominating background from inclusive neutron photoproduction makes it highly unlikely this process can be measured at the EIC without dedicated instrumentation. If one is optimistic, a dedicated hermetic veto detector may be able to deliver the necessary background suppression, depending on the final implementation of the far-forward region and the level of machine backgrounds. Such a veto system could be incorporated into the far-forward region of a second EIC detector.
@@ Acknowledgments
I would like to acknowledge Sylvester Joosten, Jihee Kim, Noémie Pilleux, and Maria Żurek for useful discussions. I also thank Lindsay DeWitt for editing this manuscript. This work was supported by the U.S. Department of Energy under Contract No. DE-AC02-06CH11357.
@@
This section provides more detail on how the rates for leading neutron production were estimated based on the HERA data of Refs. Chekanov et al. (2002, 2007). The cross sections reported by ZEUS are $\gamma p$ cross sections and therefore must have the photon flux included to obtain an $ep$ cross section. The photon flux from the electron as a function of $Q^{2}$ and $y$ is

### v2

@@ IV Veto System
As an example of a successful veto system, layered plastic scintillator detectors with single-layer efficiencies of 99.8% for single charged particles were built and operated for the NA62 experiment Ambrosino:2015ruf. Three layers of similar detectors arranged continuously around the beampipe in a configuration similar to the one sketched in Fig. 6 could provide the necessary veto efficiency of $\approx 10^{6}$. The shape of the scintillators will necessarily conform to the final shape of the beampipe, which has a complicated geometry in the far-forward region Aschenauer:2025mku.
@@
A unique challenge for a scintillator-based charged particle veto system at the EIC is that the instantaneous rate of hits per channel can be large, depending on the granularity that is chosen. The expected averaged event rate at the EIC at top luminosity is around 500 kHz, and a large fraction of those events will produce hits in the forward region. A non-negligible hit rate from beam-gas events and synchrotron radiation can also be expected. The bunches at the EIC are spaced by roughly 10 nanoseconds. Two closely timed hits, e.g. from two consecutive bunch crossings, in the scintillator can result in the later hit not being properly vetoed if the response time of the scintillator and photosensor is too long. From that perspective, a photosensor with a fast rise- and fall-time such as an MCP-PMT may be desired to avoid hits shortly after one another piling up and defeating the veto. The decay time of plastic scintillator, for example formulas based on PPO fluorescence with POPOP wavelength shifter, is commonly as short at 2 ns, making it ideal for this use case Eljen_EJ200_204_208_212. The expected radiation dose in the EIC far-forward region is approximately $3\times 10^{11}$ 1 MeV neutron equivalent of non-ionizing radiation and 100 krad of ionizing radiation in 100 fb-1. Plastic scintillators have been shown to retain their performance at significantly higher radiation doses Kharzheev:2019tfk. Finally, plastic scintillator can be extruded or molded into awkward shapes, as will likely be necessary for hermetic coverage around the beampipe.
V Projections
Figure 7: Left: Projected yields as a function of $|t|$ for 500 fb-1. Center: Background-subtracted CCE cross section. Right: Projected target-spin asymmetry including uncertainty on the background asymmetry $A_{bg}$. See text for more details.
Assuming a veto system capable of reducing the number of background events to 5000 per 100 fb-1, we make projections for the measurement of the CCE cross section and longitudinal target-spin asymmetry in Fig. 7. The shape of the background is assumed to have the same functional form as the signal shown in Fig. 4, but with a different dipole mass of 0.8 GeV to avoid bias from having exactly the same background and signal distribution shapes. An integrated luminosity of 500 fb-1, corresponding to 5 years of running at the design luminosity, is used for these projections.
Projections for the measurement of the longitudinal target-spin asymmetry $A_{UL}$ are shown in the right panel of Fig. 7. A 70% polarization on the proton beam is assumed; the impact of the nominal 1% uncertainty on the value of the polarization Huang:2025gqx; Rathmann:2025jgp is negligible compared to the statistical uncertainty. The longitudinal target-spin asymmetry of the background process $e\vec{p}\rightarrow n+X$ is currently unknown, but its magnitude will impact how well the CCE asymmetry can be extracted. To account for a lack of knowledge in the asymmetry of the background, we assume the background has an asymmetry of $0.05\pm 0.025$ and propagate this uncertainty into the final uncertainty on the measured CCE asymmetry. The resulting uncertainty is shown in the right panel of Fig. 7. This source of uncertainty can be reduced significantly using even low-luminosity data from the early runs of the EIC to measure the $e\vec{p}\rightarrow n+X$ asymmetry. The datapoints are scattered around a nominal asymmetry of $A_{UL}=1$ according to their uncertainties to mimic a real measurement. The results on $A_{UT}$ would be the similar to those shown in Fig. 7, barring a very large $A_{UT}$ for the background process $e\vec{p}\rightarrow n+X$.
The results of Ref. Tomalak:2020zlv show that a variation of $F_{A}=(1\pm 0.2)\times F_{A}^{\mathrm{Exp.}}$, where $F_{A}^{\mathrm{Exp.}}$ is the nominal experimental value of $F_{A}$, produces a roughly 10% change in $A_{UL}$ for $|t|>0.5\mathrm{~GeV^{2}}$. Below $|t|=0.5\mathrm{~GeV^{2}}$ the asymmetry depends more weakly on $F_{A}$. Therefore, it seems unlikely that this method will produce a competitive measurement of $F_{A}$, however it can be used to cross-check existing results and theoretical understanding with a novel technique [^3] One is reminded of the proton elastic form factor puzzle JeffersonLabHallA:1999epl, where the inclusion of the polarization transfer technique revealed a large discrepancy with the standard Rosenbluth method..
VI Remarks on Charged Current Cross Sections
@@
It can be seen from Fig. 2 that the CCE total cross section remains approximately constant with $\sqrt{s}$ due to competition between the nucleon form factors and increased effective flux of $W^{-}$. On the other hand, the inclusive charged-current cross section increases linearly with $E_{\nu}$ due to increasing contribution of charged-current DIS, which dominates above $E_{\nu}\approx 20$ GeV. The phase space in $\gamma p$ center-of-mass energy $W$ for hard exclusive processes such as deeply virtual Compton scattering (DVCS) and deeply virtual meson production (DVMP) are much larger than for elastic scattering, and the DVCS and DVMP cross sections are known to grow with $W$ H1:2005gdw; ZEUS:2008hcd; H1:2009cml; Kumericki:2009uq. It may be the case then that at EIC energies, the cross sections for charged-current exclusive processes such as CCDVCS ($e^{-}p\rightarrow\nu_{e}n\gamma$) or CCDVMP (e.g. $e^{-}p\rightarrow\nu_{e}p\pi^{-}$) could be larger than the CCE cross section and counterintuitively may be more experimentally favorable. In particular, the $e^{-}p\rightarrow\nu_{e}p\pi^{-}$ reaction would have improved resolution on missing transverse momentum and missing mass since they could be measured in trackers. The charged-current pion production process is also expected to have large spin asymmetries that can be exploited Graczyk:2021oyl; Graczyk:2023lrm. Theoretical estimates of these cross sections at EIC energies would therefore be interesting to see.
VII Conclusion
@@
The charged-current elastic scattering process provides a novel means to study the spin structure of the nucleon and potentially some presently unexplored corners of BSM physics. With its high luminosity and polarized beams of electrons and protons, the EIC is, in principle, an excellent tool for studying this reaction. The cross section for CCE scattering at the EIC is expected to be around 3.5 fb assuming an equal mix of left- and right-handed electrons in the beam. Target-spin asymmetries are predicted to be on the order of 100%, with their magnitude being sensitive to the axial form factor of the nucleon. However, the presence of a dominating background from inclusive neutron photoproduction makes it highly unlikely this process can be measured at the EIC without dedicated instrumentation. If one is optimistic, a dedicated hermetic veto detector may be able to deliver the necessary background suppression, depending on the final implementation of the far-forward region and the level of machine backgrounds. Such a veto system could be incorporated into the far-forward region of a second EIC detector. Even with such a veto system, it is unlikely that the EIC will deliver a competitive measurement of the axial form-factor, however the potential exists to measure target-spin asymmetries of this process for the first time, cross-checking the expectations from theory.
@@ Acknowledgments
I would like to acknowledge Tim Hobbs, Garth Huber, Sylvester Joosten, Jihee Kim, Zein-Eddine Meziani, Noémie Pilleux, and Maria Żurek for useful discussions and comments. I also thank Lindsay DeWitt for editing this manuscript. This work was supported by the U.S. Department of Energy under Contract No. DE-AC02-06CH11357.
@@
This section provides more detail on how the rates for leading neutron production were estimated based on the HERA data of Refs. ZEUS:2002gig; ZEUS:2007knd. The cross sections reported by ZEUS are $\gamma p$ cross sections and therefore must have the photon flux included to obtain an $ep$ cross section. The photon flux from the electron as a function of $Q^{2}$ and $y$ is